A game coniss of elecuna a hreedi gitube ber ucu ccte paid you te. Eadl. uess 2 $700 a -Hexes eneeuwinning number, if a ferson r one Vin cau ae erpe or to
Answer #1

Probability = Favorable Outcomes/Total Outcomes

Favorable outcomes = 1 (There is only 1 number which will be correct)

Total Outcomes = 10 * 10 * 10 = 1000

Probability of getting the 3 digit number correct = 1/1000

Probability of not getting the correct number = 1 - 1/1000 = 999/1000

Net Amount won if the correct number comes = 700 - 1 = $699

Net amount lost if correct number does not come = $1

Therefore Expected Value per day = SUM[x * P(x)] = [699 * 1/1000] - 1 * 999/1000 = -300/1000 = -3/10 = -0.3 (or $0.3 loss)

Therefore if the game is played for a year, expected loss = -0.3 * 365 = -$109.5   ($109.50 loss)